OpenStudy (anonymous):

A pet shop had 8 identical fish tanks in the shape of rectangular prisms. The width and height of each tank was (x - 7) inches and the length was 8 inches more than its width. The expression below shows the maximum amount of water which all the tanks together can hold. 8(x + 1)(x - 7)to the 2nd power What does the factor (x + 1)(x – 7)to the 2nd power represent?

OpenStudy (anonymous):

Help!

OpenStudy (ranga):

Volume = L * W * H W = (x-7) H = (x-7) L = W + 8 = x - 7 + 8 = x + 1 L = (x+1) Volume of one fish tank = L * W * H = (x+1) * (x-7) * (x-7) = (x+1)(x-7)^2